# Theory of two-component superfluidity of microcavity polaritons

### v1

@@ Abstract
We investigate Bose–Einstein condensation and two-component superfluidity composed of upper (UP) and lower (LP) polaritons confined to a two-dimensional microcavity. By studying a modified Hamiltonian which incorporates interactions between both polariton branches, we derive the collective excitation spectrum and express the sound velocity as a function of detuning and Rabi splitting. Our analysis reveals that the interplay between effective masses and exciton–photon fractions in the UP and LP branches can significantly enhance the superfluid properties compared to a one-component condensate of only lower polaritons. Specifically, we discover that larger Rabi splitting increases the sound velocity and raises the critical temperature for superfluidity, while negative and positive detuning further favors a two-component condensate. Additionally, we demonstrate that these trends are consistent and applicable to various materials, including GaAs and the transition metal dichalcogenides (TMDCs), thereby indicating the potential for high-temperature superfluid phenomena in two-dimensional polariton systems. We hope these findings will provide a deeper understanding of the physics of structures containing multi-component polariton systems and generate experimental realizations of tunable quantum fluids with light.
@@ I Introduction
Bose-Einstein condensation (BEC) and superfluidity in exciton-polariton systems have been emerging as profoundly interesting areas of investigation in condensed matter physics. This is due to the unique properties of polaritons which are quasiparticles arising from the strong coupling between excitons and photons Moskalenko2000 ; Littlewood2007 . These hybrid particles exhibit extremely light effective masses, leading to high critical temperatures for condensation. The critical temperature is relatively high due to the inverse relationship between the 2D thermal de Broglie wavelength and the mass of the quasiparticle. As this wavelength approaches the interparticle spacing, Bose-Einstein condensation (BEC) and superfluidity of exciton-polaritons have been experimentally observed in microcavities Carusotto2013 ; Proukakis2017 ; Peng2022 . Much of the existing research focuses on polaritons in microcavities and under external potentials such as harmonic and anharmonic traps berman2008 .
When a photon is absorbed in a semiconductor an electron is excited from the valence to the conduction band, leaving behind a positively charged vacancy or “hole”. The electron–hole pair can be bound by the Coulomb attraction, thereby forming a quasiparticle which is referred to as an exciton exciton1983 . Under certain densities, these excitons are predicted to exhibit BEC and superfluidity at temperatures significantly higher than those achievable with atomic gases butov2003 . Exciton-polaritons, formed when these excitons strongly couple with photons in an optical microcavity, exhibit both light- and matter-like characteristics deng2010 ; snokeKeeling2017 . This unusual superposition facilitates macroscopic quantum states at relatively high temperatures, making polaritons an attractive means for investigating BEC and superfluidity in two-dimensional (2D) systems Carusotto2013 ; Proukakis2017 ; Peng2022 . Studies have also shown that exciton superfluidity can remain robust even with disorder or magnetic field effects berman2006 . In addition to condensation and superfluidity, polaritons can undergo parametric scattering and exhibit optically parametrized oscillations ciuti2003 .
Previous work on polariton condensation has demonstrated the existence of coherent states in semiconductor microcavities. This has been accompanied by experimental observations in GaAs berman2008 and CdTe at cryogenic temperatures richard2005 . Recent advances in materials science and engineering have expanded these investigations to include monolayer transition metal dichalcogenides (TMDCs) manzeli2017 and organic materials lerario2017 , expanding the realms of polariton condensation.
Two-dimensional van der Waals materials, such as atomically thin transition metal dichalcogenides, possess very distinctive physical properties which make them appealing for a wide range of applications. Monolayers of TMDCs such as MoS2, MoSe2, MoTe2, WS2, WSe2, and WTe2 function as 2D direct-bandgap semiconductors with various applications in electronics and optoelectronics mak2010 . Interest in these monolayers stems from their direct bandgap structure splendiani2010 , the existence of excitonic valley physics, and the feasibility of electrically tunable, strong light–matter interactions xiao2012 . Monolayer TMDCs have already been adopted in field-effect transistors, logic devices, and in both lateral and tunneling optoelectronic designs wang2012 . They exhibit hexagonal lattice structures, with the K and K points in the hexagonal Brillouin zone representing the valleys in the valence and conduction bands Kormanyos2015 . Their excitonic properties have been investigated in numerous experimental and theoretical studies xu2014 . The notable binding energy and extended lifetime of interlayer excitons in van der Waals heterostructures rivera2015 have also motivated substantial research hong2014 . Exciton–polaritons in a TMDC monolayer inside a microcavity were observed experimentally at room temperatures liu2015 , while the superfluidity of exciton-polaritons in such TMDC-based microcavities has been examined in Refs dufferwiel2015 ; kogar2017 . Doping can result in the creation of trions in MoS2, which possess significant binding energies Mak2013 . Furthermore, the ability to create heterostructures by stacking TMDC layers enables the formation of interlayer excitons, which exhibit long lifetimes and robust coherence properties rivera2015 . TMDC-based polaritonic systems embedded in microcavities have shown superfluidity at room temperature, underscoring their potential for future quantum technologies Peng2022 . Besides TMDC double layers, dipolar excitons can also achieve superfluidity in more exotic lattices such as $\alpha$-T3, under appropriate mass and coupling conditions berman2022nanomaterials Exotic phases such as polaritonic or excitonic time crystals have been proposed in TMDC strips subjected to periodic potentials martins2023 , highlighting the breadth of many-body phenomena accessible in 2D materials. Exciton-related phenomena are not limited to conventional semiconductors and TMDCs. Recent studies extend to topological nodal-line semimetals, revealing unique anisotropic optical responses and potential excitonic states in exotic band structures kandel2023 ; kandel2024 .
Gallium Arsenide (or GaAs) is a widely studied III–V semiconductor with a direct bandgap, making it highly suitable for microcavity-based optoelectronic and photonic applications weisbuch1992 . Specifically, GaAs quantum wells embedded within optical microcavities enable the realization of strong exciton-photon coupling and the formation of exciton-polaritons Skol1998 . This strong coupling gives rise to a characteristic Rabi splitting in the excitation spectrum, which underpins the observation of polaritonic effects such as condensation and superfluid-like behavior at comparatively high temperatures Sanvitto2016 . The well-developed epitaxial growth techniques for GaAs allow for precise control over layer thickness, doping profiles, and defect density, ensuring high-quality microcavity structures with minimal disorder steger2015 . These properties, alongside the large oscillator strength and relatively narrow excitonic linewidths in GaAs quantum wells khitrova2006 favor low-threshold polariton lasing zhao2021 and stable Bose–Einstein condensation of polaritons kasprzak2006 . GaAs-based devices serve as a versatile platform for investigating fundamental light–matter interactions, as well as for designing novel photonic components such as polariton lasers, optical switches, and other polariton-based quantum technologies.
Exciton-polaritons (henceforth referred to simply as “polaritons”) form in semiconductor microcavities through the strong coupling of quantum well excitons with photons confined between two Bragg mirrors szymanska2004 . The coupling strength is manifested as Rabi splitting in the excitation spectrum, resulting in two distinct branches: the lower polariton (LP) and the upper polariton (UP). The LP branch has a lower energy and an extremely small effective mass, allowing it to behave as a two-dimensional weakly interacting Bose gas. Experiments have shown that polaritons can exhibit both BEC and superfluidity in microcavity systems snokecoherent2006 . Their tunable interactions, achieved through external fields or cavity-parameter adjustments, render polaritons ideal for studying quantum many-body phenomena and for realizing next-generation photonic devices.
Recent experimental observations Chen2023 have demonstrated Bose–Einstein condensation of upper polaritons in a WS2-based microcavity, showing that under specific detuning and pumping conditions, condensation in the upper polariton branch can become the dominant phase. Additionally, experiments Comaron2025 have highlighted how tunable polariton interactions in driven-dissipative systems enhance spatial coherence and enable Berezinskii–Kosterlitz–Thouless-like crossovers, consistent with interaction-driven superfluid behavior. These developments underscore the importance of incorporating both upper and lower polariton condensates in a unified theoretical description of polariton superfluidity and Bose–Einstein condensation.
In this paper, we direct our attention to a simpler yet equally rich scenario - a two-component condensate of both upper and lower polaritons without the influence of in-plane potentials. We allow scattering between the upper and lower polariton branches, and investigate how this coupling affects the superfluid transition and the resulting collective excitations. By adapting approaches previously applied to two-component exciton superfluidity arafat2024semi , our study provides new insights into multi-component polariton system’s behavior and clarifies how superfluidity can emerge under simpler physical conditions. We further analyze the way in which detuning plays a role in determining the critical temperature and sound velocity, thereby shedding light on the role of multi-component dynamics in polaritonic superfluidity. We find that the Bose gas of polaritons, forming a two-component Bose-Einstein condensate of lower and upper polariton, demonstrate a higher sound velocity and critical temperature rather than just a one-component condensate of lower polaritons. These findings build upon recent advances in polariton condensates and highlight their promise for controlling light-matter interactions in future quantum technologies Sanvitto2016 ; alyatkin2020 ; Bhattacharya2013 . In this work, we show that the effective masses of the upper and lower polariton branches, along with their exciton and photon fractions, vary significantly with detuning, leading to a richer interplay in the two-component condensate than in a one-component condensate. We find that the sound velocity in the two-component condensate is generally higher than in the single-component condensate, and that the critical temperature for superfluidity is also enhanced across all detuning, when other parameters are held constant. Increasing the Rabi splitting consistently raises the sound velocity and therefore the critical temperature, underscoring the benefits of stronger light–matter coupling. These trends hold in both GaAs and various TMDC materials examined, indicating their potential as versatile candidate materials for polariton superfluidity at higher temperatures. Our two-component approach offers enhanced superfluid properties through additional interaction channels and opens new avenues for parameterized engineering in polaritonic devices, potentially enabling high-temperature operation.
The rest of the paper is organized as follows. In Section II, we modify the Hamiltonian for a two-dimensional system of excitons for the Bose gas of polaritons, forming a two-component Bose-Einstein condensate of lower and upper polaritons. We reexpress it in the basis of upper and lower polariton operators, carefully ensuring that we keep the relevant interactions. Details of this process are presented in Appendix B-D. In Section III we derive the spectrum of collective excitations and carefully incorporate the Hopfield coefficients into the spectrum. In Section IV, we build on that and express the sound velocity of the two-component condensate of upper and lower polaritons as a function of these Hopfield coefficients, which allows us to investigate the manner in which the detuning and Rabi splitting affect the sound velocity. In this section we also include our comparisons with the previously investigated one-component picture of solely lower polaritons in GaAs berman2008 . In Section V we develop the subsequent mean-field model of superfluidity and find the critical temperature for superfluidity in the two-component Bose gas of upper and lower polaritons as a function of various parameters. In Section VI, we further our analysis by using various TMDCs each with their own effective masses of excitons. We also compare the two-component condensate and superfluid to a model one-component condensate and superfluid in various TMDCs. In Section VII we present our conclusions and a discussion of our results.
@@ II Effective Hamiltonian of upper and lower polaritons
Polaritons are quantum superpositions of excitons and photons. The total Hamiltonian of excitons and microcavity photons is given by $\hat{H}_{total}~{}=~{}\hat{H}_{exc}+\hat{H}_{ph}+\hat{H}_{exc-ph}$, where $\hat{H}_{exc}$ is the excitonic Hamiltonian, $\hat{H}_{ph}$ is the photonic Hamiltonian, and $\hat{H}_{exc-ph}$ is a Hamiltonian of exciton-photon interaction. We can define each of these explicitly.
@@
where $A$ is a normalization area. Here, the first term represents the energy of the excitons, and the second term represents the exciton-exciton interaction. The operator $\hat{b}_{P}$ is the excitonic annihilation operator, and the operator $\hat{b}^{\dagger}_{P}$ is the excitonic creation operator for momentum $P$. The contact potential $U_{q}$ is approximated by Eq. (12). The exciton energy dispersion $\varepsilon_{\text{ex}}(P)$ is given by
@@
where $\hbar\Omega_{R}$ is the Rabi splitting, representing the coupling strength between photons and excitons.To fully describe the system, we express the Hamiltonian in the basis of upper and lower polariton operators. We substitute the following expressions from Equation (6) of Ref. hopfield1958 .
@@
$\hat{H}_{0}=\sum_{\textbf{P}}\varepsilon_{LP}(P)\hat{l}_{\textbf{p}}^{\dagger}\hat{l}_{\textbf{p}}+\sum_{\textbf{P}}\varepsilon_{UP}(P)\hat{u}_{\textbf{p}}^{\dagger}\hat{u}_{\textbf{p}}$ (11)
@@
where, the contact potential $U_{q}$ is approximated by
$U_{q}=\frac{6e^{2}a_{2D}}{\epsilon}$ (12)
and $a_{2D}$ is the effective 2D Bohr radius of excitons,
@@
This becomes
$\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{q}\left(X_{P+q}\hat{l}_{P+q}^{\dagger}-C_{P+q}\hat{u}_{P+q}^{\dagger}\right)\left(X_{P^{\prime}-q}\hat{l}_{P^{\prime}-q}^{\dagger}-C_{P^{\prime}-q}\hat{u}_{P^{\prime}-q}^{\dagger}\right)\left(X_{P}\hat{l}_{P}-C_{P}\hat{u}_{P}\right)\left(X_{P^{\prime}}\hat{l}_{P^{\prime}}-C_{P^{\prime}}\hat{u}_{P^{\prime}}\right)$ (14)
@@
$\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{q}\left(X_{P+q}\hat{l}_{P+q}^{\dagger}-C_{P+q}\hat{u}_{P+q}^{\dagger}\right)\left(X_{P^{\prime}-q}\hat{l}_{P^{\prime}-q}^{\dagger}-C_{P^{\prime}-q}\hat{u}_{P^{\prime}-q}^{\dagger}\right)\left(X_{P}\hat{l}_{P}-C_{P}\hat{u}_{P}\right)\left(X_{P^{\prime}}\hat{l}_{P^{\prime}}-C_{P^{\prime}}\hat{u}_{P^{\prime}}\right)$
@@
Certain terms in the expanded Hamiltonian represent interactions that are not typical for polaritons due to the following reasons. Polaritons in the upper ($\hat{u}$) and lower ($\hat{l}$) branches generally do not transform into one another without a strong coupling mechanism, especially with the large Rabi splitting. The Rabi splitting represents the energy difference between the upper and lower polariton branches, and when this splitting is significant (on the order of tens of meV), transitions between these branches require a comparable amount of energy, making spontaneous interbranch conversions highly improbable Peng2022 ; Carusotto2013 . As a result, terms where a lower and an upper polariton combine to form two of the same type (either both lower or both upper) are not physically consistent. Processes that would convert one lower and one upper polariton into two lower or two upper polaritons are unlikely to occur without the input of additional energy. This makes certain cross-type conversion terms energetically unfavorable, if even possible, so we discard them from consideration. After removing these non-physical terms, the interaction Hamiltonian simplifies to
$\hat{H}_{\text{int}}=\frac{1}{2A}\sum_{\begin{subarray}{c}\mathbf{P},\mathbf{P}^{\prime},\mathbf{q}\end{subarray}}U_{q}\cdot\frac{1}{4}\Big{[}$ $\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{l}_{\mathbf{P}^{\prime}}+\hat{u}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}$ (15)
$+\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}+\hat{u}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}}\hat{l}_{\mathbf{P}^{\prime}}$
$+\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}}\hat{l}_{\mathbf{P}^{\prime}}+\hat{u}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}\Big{]}$

### v2

@@ Abstract
We develop a microscopic mean-field theory describing the coexistence of Bose–Einstein condensates of upper and lower polaritons (UP/LP) in a semiconductor microcavity. Incorporating interbranch scattering within a modified polariton Hamiltonian, we introduce a phenomenological population–split parameter $\alpha$ that quantifies the relative LP/UP occupations. At zero detuning, the critical temperature becomes independent of $\alpha$, converging to a single value that marks the balanced, resonant regime. Away from resonance, variations in $\alpha$ lead to distinctive and experimentally resolvable changes in both the sound velocity $c_{s}$ and critical temperature $T_{c}$, relative to the single-component (LP-only) condensate limit. The system under study consists of excitons confined in A transition metal dichalcogenide (TMDC) monolayer, particularly WSe2 embedded within a planar optical microcavity of GaAs where they strongly couple to cavity photons. Our analysis focuses on monolayer WSe2 embdedded in a GaAs microcavity. We present results for GaAs/AlGaAs quantum wells embedded in a GaAs microcavity in the Appendix.While mean-field in scope, the framework provides analytic benchmarks and physical insight for future treatments that include dissipation and fluctuations in nonequilibrium polariton superfluids.
@@ I Introduction
Bose–Einstein condensation (BEC) and superfluidity of exciton-polaritons have become key topics in two-dimensional (2D) systems due to the hybrid light–matter nature of polaritons deng2010 ; snokeKeeling2017 ; kavokin2017 , formed via strong coupling between quantum well excitons and cavity photons Moskalenko2000 ; Littlewood2007 . Excitons are bound electron–hole pairs that form in semiconductors through Coulomb attraction. When such excitons couple strongly to the quantized electromagnetic field of an optical microcavity, they hybridize with cavity photons to form new light–matter quasiparticles known as exciton–polaritons weisbuch1992 . In GaAs/AlGaAs structures, the excitons are confined within a quantum well embedded between distributed Bragg reflectors that define the optical cavity Skol1998 . In transition metal dichalcogenide (TMDC) realizations, by contrast, the excitons reside in an atomically thin monolayer placed inside a similar cavity, where the large binding energy enhances the strength of light–matter coupling dufferwiel2015 . Their low effective mass enables high critical temperatures for superfluidity  Carusotto2013 ; Proukakis2017 ; Peng2022 , with robust behavior under disorder and external fields berman2006 . Polaritons support parametric scattering ciuti2003 ; savona2005 and coherence across diverse materials, including GaAs berman2008 , CdTe richard2005 , and TMDCs manzeli2017 ; lerario2017 , where Bose-Einstein condensation kasprzak2006 and room-temperature superfluidity have been observed dufferwiel2015 ; kogar2017 ; Peng2022 ; zhao2021 , alongside promising optoelectronic applications mak2010 and the emergence of topological excitations in spinor condensates rubo2007 . Theoretical models have further explored nonlinear polariton condensation beyond the linear regime eastham2001 . Additionally, experiments Comaron2025 have highlighted how tunable polariton interactions in driven-dissipative systems enhance spatial coherence and enable Berezinskii–Kosterlitz–Thouless-like crossovers, consistent with interaction-driven superfluid behavior.
The study of polariton condensates naturally extends to a broader class of multi-component quantum fluids that exhibit rich collective phenomena. Multi-component quantum condensates (comprising coexisting, interacting macroscopic quantum states) exhibit collective phenomena absent in single-component superfluids. In spinor atomic Bose–Einstein condensates, simultaneous occupation of multiple hyperfine states gives rise to coupled spin-density modes, relative-phase dynamics, and topological defects such as half-quantum vortices Ho1998 ; Stamper2013 . In magnon systems, multi-branch condensation produces hybridized magnon modes and interaction-mediated phase transitions Demokritov2006 ; Aeppli2022 .An instructive parallel can be drawn with superfluid 3He, where distinct pairing symmetries (A and B phases) correspond to multiple order parameters with coupled collective excitations Vollhardt2003 ; Leggett1975 . Although the microscopic mechanisms differ, the analogy underscores how coexisting condensate components can support richer excitation spectra and phase dynamics than their single-component counterparts.
Microcavity polaritons offer a uniquely tunable platform for exploring two-component condensate physics. The upper (UP) and lower (LP) polariton branches arise naturally from strong exciton–photon coupling, with an energy splitting that can be engineered through cavity design. These two branches correspond to the eigenmodes of the coupled exciton–photon system obtained through the Hopfield diagonalization. The lower polariton (LP) has a larger excitonic fraction near positive detuning, giving it a heavier effective mass and stronger interactions, whereas the upper polariton (UP) becomes more photonic at negative detuning and has a lighter mass. The two modes are separated by the vacuum Rabi splitting $\hbar\Omega_{R}$, which quantifies the strength of the light–matter coupling and sets the energy scale for the polariton system Carusotto2013 ; weisbuch1992 . While most experiments observe condensation exclusively in the lower branch due to efficient phonon-assisted relaxation kasprzak2006 , recent experimental work has demonstrated upper-branch condensation Chen2023 , suggesting that controlled dual-branch occupation may be achievable under appropriate conditions. A key control parameter in such systems is the exciton–photon detuning, $\Delta_{0}$, defined as the energy difference between the bare cavity photon mode and the exciton resonance. Detuning can be tuned in situ by adjusting the cavity length, refractive index (via temperature or angle of incidence), or through fabrication of distributed Bragg reflectors of different thickness dufferwiel2015 . Varying $\Delta_{0}$ changes the Hopfield coefficients that determine the excitonic and photonic fractions of the polariton branches, thereby modifying their effective masses and interaction strengths. Near zero detuning, both components are equally mixed, while large positive or negative detuning yields predominantly excitonic or photonic character, respectively. This tunability, along with control over excitation density and bare Rabi splitting, allows polaritons to access regimes ranging from light-mass, weakly interacting fluids to heavy-mass, strongly interacting excitonic gases.
Despite these experimental advances and the theoretical progress in understanding single-branch polariton condensates, no microscopic theory has yet incorporated mutual coherence and interaction between lower and upper polariton (LP and UP) condensates in a unified superfluid framework. The upper branch, despite experimental signatures of occupation Chen2023 , remains largely absent from such treatments. Previous work has focused largely on one-component polariton condensates, particularly in the lower polariton (LP) branch. Polariton condensates can form in two distinct branches—lower polariton (LP) and upper polariton (UP)—arising from Rabi splitting szymanska2004 ; snokecoherent2006 .
Motivated by advances in multi-component superfluidity and tunable quasiparticle engineering arafat2024semi ; Sanvitto2016 ; alyatkin2020 ; Bhattacharya2013 , we develop a theory of superfluidity in a two-component polariton condensate comprising both LP and UP branches, formulated within an equilibrium mean-field (Bogoliubov) framework that explicitly includes detuning dependence and interaction-driven effects. Our analysis is carried out in the long-wavelength limit, where the lower and upper polariton dispersions are expanded to quadratic order in momentum, and the effective masses treated as constants, allowing analytical access to the collective modes and superfluid parameters such as the concentration of the superfluid component and the critical temperature for superfluidity. Our theoretical framework treats the condensate population distribution between LP and UP branches as a phenomenological input, characterized by a parameter $\alpha\in[0,1]$, and calculates the resulting superfluid properties via mean-field theory. The systems we consider in particular are GaAs/AlGaAs quantum wells embedded in a GaAs microcavity, and TMDC (WSe2) monolayers embedded in a GaAs microcavity regan2022 . Our work establishes theoretical benchmarks for an experimentally relevant regime, soon accessible with current advances in polariton microcavities. We provide quantitative predictions for the collective excitation spectrum, sound velocity, and critical temperature for superfluidity that would uniquely distinguish genuine two-component LP–UP superfluidity from single-branch superfluidity, thereby offering clear signatures for future experimental verification.
[…]
We show that this two-component structure yields enhanced and tunable sound velocity and critical temperature of superfluidity compared to the LP-only case. Our results are general, material-independent, and analytically tractable—offering a predictive framework for future studies of nonequilibrium polariton dynamics, Lindblad-driven condensates, and Josephson-like phase control in strongly coupled light–matter systems.
The remainder of this paper is organized as follows. Section II introduces the modified Hamiltonian for upper and lower polaritons, including the systematic reduction of interaction terms. Section III derives the collective excitation spectrum within the Bogoliubov framework. Section IV examines the tunable sound velocity and its dependence on detuning, Rabi splitting, and the population–split parameter $\alpha$. Section V develops the superfluid criterion and derives analytic expressions for the critical temperature for superfluidity and concentration. In this section we also analyze how the critical temperature for superfluidity depends on the detuning, Rabi splitting, and population-split parameter $\alpha$, and demonstrate a retrieval procedure that allows one to infer $\alpha$ from measured values of the critical temperature $T_{c}$ and critical density $n_{c}$. Section VI concludes with a discussion of the main findings and their implications for tunable two-component polariton superfluidity. Appendices A and B present detailed derivations of the interaction reduction and polariton effective masses that are used in our calculations. Appendix C presents all pertinent results of our analysis for GaAs/AlGaAs embedded in GaAs microcavities.
@@ II Effective Hamiltonian of upper and lower polaritons
Polaritons are quantum superpositions of excitons and photons. The total Hamiltonian of excitons and microcavity photons is given by $\hat{H}_{\text{total}}~=~\hat{H}_{\text{exc}}+\hat{H}_{\text{ph}}+\hat{H}_{\text{exc-ph}}$, where $\hat{H}_{\text{exc}}$ is the excitonic Hamiltonian, $\hat{H}_{\text{ph}}$ is the photonic Hamiltonian, and $\hat{H}_{\text{exc-ph}}$ is a Hamiltonian of exciton-photon interaction. We can define each of these explicitly.
@@
where $A$ is a normalization area. Here, the first term represents the energy of the excitons, and the second term represents the exciton-exciton interaction. The operator $\hat{b}_{P}$ is the excitonic annihilation operator, and the operator $\hat{b}^{\dagger}_{P}$ is the excitonic creation operator for momentum $P$. The contact potential $U_{q}$ is approximated by Eq. (12) and from experimental benchmarks. The exciton energy dispersion $\varepsilon_{\text{ex}}(P)$ is given by
@@
where $\hbar\Omega_{R}$ is the Rabi splitting, representing the coupling strength between photons and excitons.To fully describe the system, we express the Hamiltonian in the basis of upper and lower polariton operators. We substitute the following expressions from Eq. (6) of Ref. hopfield1958 .
@@
$\hat{H}_{0}=\sum_{P}\varepsilon_{LP}(P)\hat{l}_{P}^{\dagger}\hat{l}_{P}+\sum_{P}\varepsilon_{UP}(P)\hat{u}_{P}^{\dagger}\hat{u}_{P}$ (11)
@@
where, the contact potential $U_{q}$ for the exciton-exciton interaction is approximated to $U_{0}$, and is defined for GaAs/AlGaAs as Carusotto2013
$U_{0}=\frac{6e^{2}a_{2D}}{\epsilon}$ (12)
and $a_{2D}$ is the effective 2D Bohr radius of excitons.
@@
It is crucial to note that for monolayer WSe2, the exciton–exciton interaction is more accurately described by the Rytova-Keldysh potential Berkelbach2013 , which accounts for reduced screening in atomically thin materials. Rather than deriving $U_{0}$ microscopically, we adopt the experimentally measured value from polariton blueshift measurements Kravtsov2020 ; flatten2016 . Proceeding we get the following
$\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{0}\left(X_{P+q}\hat{l}_{P+q}^{\dagger}-C_{P+q}\hat{u}_{P+q}^{\dagger}\right)\left(X_{P^{\prime}-q}\hat{l}_{P^{\prime}-q}^{\dagger}-C_{P^{\prime}-q}\hat{u}_{P^{\prime}-q}^{\dagger}\right)\left(X_{P}\hat{l}_{P}-C_{P}\hat{u}_{P}\right)\left(X_{P^{\prime}}\hat{l}_{P^{\prime}}-C_{P^{\prime}}\hat{u}_{P^{\prime}}\right)$ (14)
@@
$\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{0}\left(X_{P+q}\hat{l}_{P+q}^{\dagger}-C_{P+q}\hat{u}_{P+q}^{\dagger}\right)\left(X_{P^{\prime}-q}\hat{l}_{P^{\prime}-q}^{\dagger}-C_{P^{\prime}-q}\hat{u}_{P^{\prime}-q}^{\dagger}\right)\left(X_{P}\hat{l}_{P}-C_{P}\hat{u}_{P}\right)\left(X_{P^{\prime}}\hat{l}_{P^{\prime}}-C_{P^{\prime}}\hat{u}_{P^{\prime}}\right)$
@@
Certain terms in the expanded Hamiltonian represent interactions that are not typical for polaritons due to the following reasons. Polaritons in the upper ($\hat{u}$) and lower ($\hat{l}$) branches generally do not transform into one another without a strong coupling mechanism, especially with the large Rabi splitting. The Rabi splitting represents the energy difference between the upper and lower polariton branches, and when this splitting is significant (on the order of tens of meV), transitions between these branches require a comparable amount of energy, making spontaneous interbranch conversions highly improbable Peng2022 ; Carusotto2013 . As a result, terms where a lower and an upper polariton combine to form two of the same type (either both lower or both upper) are not physically consistent and are sorted through in Appendix A. Processes that would convert one lower and one upper polariton into two lower or two upper polaritons are unlikely to occur without the input of additional energy. This makes certain cross-type conversion terms energetically unfavorable, if even possible, so we discard them from consideration. After removing these non-physical terms, the interaction Hamiltonian simplifies to
$\hat{H}_{\text{int}}=\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{0}\Big[$ $g_{\ell\ell}^{PP^{\prime}q}\,\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}+g_{uu}^{PP^{\prime}q}\,\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$ (15)
$+g_{\ell u}^{PP^{\prime}q}\,\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}+g_{u\ell}^{PP^{\prime}q}\,\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$+\tilde{g}_{\ell u}^{PP^{\prime}q}\,\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}+\tilde{g}_{u\ell}^{PP^{\prime}q}\,\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}\Big].$
With the Hopfield factor coefficients defined as
$g_{\ell\ell}^{PP^{\prime}q}$ $=X_{P+q}X_{P^{\prime}-q}X_{P}X_{P^{\prime}},$ (16)
$g_{uu}^{PP^{\prime}q}$ $=C_{P+q}C_{P^{\prime}-q}C_{P}C_{P^{\prime}},$ (17)
$g_{\ell u}^{PP^{\prime}q}$ $=X_{P+q}C_{P^{\prime}-q}X_{P}C_{P^{\prime}},$ (18)
$g_{u\ell}^{PP^{\prime}q}$ $=C_{P+q}X_{P^{\prime}-q}C_{P}X_{P^{\prime}},$ (19)
$\tilde{g}_{\ell u}^{PP^{\prime}q}$ $=X_{P+q}C_{P^{\prime}-q}C_{P}X_{P^{\prime}},$ (20)
$\tilde{g}_{u\ell}^{PP^{\prime}q}$ $=C_{P+q}X_{P^{\prime}-q}X_{P}C_{P^{\prime}}.$ (21)